OpenStudy (anonymous):

how to i do sec theta=4, and sin theta theta >0, find tan theta

OpenStudy (anonymous):

\[is~ it~\sec \theta =4~and~\sin \theta >0~,~find~\tan \theta~?\]

OpenStudy (anonymous):

\[\sin \theta(\theta)>0~? \]

OpenStudy (anonymous):

Sec α = 1/cos⁡α = 4 Sin α = 0 Tan α = sin⁡α/cos⁡α = 0/4 = 0 Is that you mean ?